# Influence of density-dependent diffusion on pattern formation in a bounded habitat

### v3

@@ V Results for space-dependent diffusivity
Typical stationary profiles of $u(x)$, $D(x)$ and the flux $J=-D(x)\partial_{x}u$ are shown in Fig. 7, changing the interface width $s$, for two different values of $D_{in}$. Increasing $s$ affects the effective width of the refuge and the wavelength of the oscillations is stretched, diminishing the number of peaks. When $s$ decreases, the intensity of the flux at ±$\pm L$ decreases, allowing the growth of the peak closer to the interface, while outside the refuge, the density becomes lower near the interface. Then in the limit $s\to 0$, a density jump is expected, as well as a jump in the first derivative of the density, such that the current $J=-D(x)\partial_{x}u$ remains continuous at the interfaces. The fluxes are maximal at the interfaces, pointing outwards. Inside the refuge, the fluxes tend to depopulate the peaks (as observed for $D_{1}$), but the logistic growth balances the fluxes and a steady state is attained. Independently of the width $s$, for $D_{in}=0.01$ (a)-(c), Eq. (9) approximately holds and oscillations are barely sustained, except for the external peak, while they are strongly damped for $D_{in}=0.02$ (d)-(f). Since small values of $s$ produce qualitatively similar results, we fix $s=0.2$.
@@
Figure 9: Phase diagrams in the plane $D_{in}-r_{in}$ for the binary profile $D(x)$. The regions I-III correspond to those described in Sec. III. The heat plots represent the real $k_{R}$ (a) and imaginary $k_{I}$ (b) parts of the dominant root of $\lambda(k)$. The symbols correspond to results of the numerical integration of Eq.-(2): oscillations are observed (filled), or not (hollow). For the plus-circles a steady state with sustained oscillations is reached. The dashed and solid lines correspond to the prediction given by Eqs. (9) and (11). Fixed parameters are $D_{out}=1$, $r_{out}=-0.02$.
@@ VI Final remarks
Acknowledgements: We are grateful to Eduardo H.F. Colombo for insightful discussions. We acknowledge partial financial support by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brazil (CAPES) - Finance Code 001. C.A. also acknowledges partial support received from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq)-Brazil (311435/2020-3) and Fundação de Amparo à Pesquisa do Estado de Rio de Janeiro (FAPERJ)-Brazil (CNE E-26/201.109/2021).

### v4

@@ V Results for space-dependent diffusivity
Typical stationary profiles of $u(x)$, $D(x)$ and the flux $J=-D(x)\partial_{x}u$ are shown in Fig. 6, changing the interface width $s$, for two different values of $D_{in}$. Increasing $s$ affects the effective width of the refuge and the wavelength of the oscillations is stretched, diminishing the number of peaks. When $s$ decreases, the intensity of the flux at ±$\pm L$ decreases, allowing the growth of the peak closer to the interface, while outside the refuge, the density becomes lower near the interface. Then in the limit $s\to 0$, a jump in the first derivative of the density is expected, such that the current $J=-D(x)\partial_{x}u$ remains continuous at the interfaces. The fluxes are maximal at the interfaces, pointing outwards. Inside the refuge, the fluxes tend to depopulate the peaks (as observed for $D_{1}$), but the logistic growth balances the fluxes and a steady state is attained. Independently of the width $s$, for $D_{in}=0.01$ (a)-(c), Eq. (9) approximately holds and oscillations are barely sustained, except for the external peak, while they are strongly damped for $D_{in}=0.02$ (d)-(f). Since small values of $s$ produce qualitatively similar results, we fix $s=0.2$.
@@
Figure 8: Phase diagrams in the plane $D_{in}-r_{in}$ for the binary profile $D(x)$. The regions I-III correspond to those described in Sec. III, delimited by dashed and solid lines given by Eqs. (9) and (12), respectively, setting $\bar{D}=D_{in}$. The heat plots represent the real $k_{R}$ (a) and imaginary $k_{I}$ (b) parts of the dominant root of $\lambda(k)$. The symbols correspond to the type of density profile observed in the numerical long-time solution of Eq. (2): sustained oscillations (plus-circles), damped oscillations (filled), and uniform state (hollow), in agreement with regions I-III. Fixed parameters are $D_{out}=1$, $r_{out}=-0.02$.
@@ VI Final remarks
Acknowledgements: We are grateful to Eduardo H.F. Colombo for insightful discussions. We acknowledge partial financial support by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brazil (CAPES) - Finance Code 001. C.A. also acknowledges partial support received from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq)-Brazil (311435/2020-3) and Fundação de Amparo à Pesquisa do Estado de Rio de Janeiro (FAPERJ)-Brazil (CNE E-26/201.109/2021,E-26/204.130/2024).